# Move and Act: Enhanced Object Manipulation and Background Integrity for Image Editing

### v1

@@ Abstract
Keywords Image Editing  $\cdot$ Text-to-Image Generation  $\cdot$ Diffusion Model
1 Introduction
Recent advancements in text-to-image generation using diffusion models [1, 2, 3, 4] have significantly improved image generation capabilities. By capitalizing on extensive training data, these models [2, 3] excel in generating remarkable images based on user prompts. Compared to generative adversarial networks (GANs), diffusion models provide superior quality and diversity in image generation. Consequently, numerous studies [5, 6, 7, 8, 9, 10] have begun investigating diffusion models for image editing tasks. Predominantly, these methods involve text-conditioned editing, where users supply text input to modify images. Although techniques like [5, 6, 7, 8] can adeptly edit image style, object appearance, and object category based on text input, they do not alter scene layout or modify object actions in the image. Nonetheless, there are instances where users desire to edit an object’s action while maintaining its appearance and background. This specific editing process is called consistent image editing.
@@ 1 Introduction
A standing cat
(e)  (a) Input
(f) (b) InstructPix2Pix
(g)  (c) PnP
(h)  (d) MasaCtrl
(i)   (e) Move&Act
Figure 1: First Row: our method allows users to control where to place the edited object. Second Row: altering only action, our method captures the “standing” pose and maintains consistent backgrounds with the input image.
Imagic [11] enables a prompt fine-tuning process for complex text-based semantic edits for consistent image editing. MasaCtrl [9] offers a tuning-free approach, transforming self-attention into mutual self-attention. These methods edit an object’s action without allowing control over the object’s generated position. DragonDiffusion [12] facilitates object displacement within the image, yet it restricts the editing of the relocated object’s action. In practical applications, users may require simultaneous editing of an object’s action and position. As illustrated in the first row of Figure 1, we allow users to input editing text to adjust the object’s action, while also providing a bounding box to dictate the object’s generated position.
We also maintain better background preservation. By designating the object’s generation position as its original location, our method effectively reduces to MasaCtrl’s editing function, solely modifying the object’s action. In the second row of Figure 1, the backgrounds generated by InstructPix2Pix [13], PnP [6] and MasaCtrl [9] exhibit a certain loss in detail, whereas our method produces a background more closely resembling the input image. Also, the cat’s action in our results aligns more accurately with the editing prompt and the surrounding environment.
@@
2 Related work
2.1 Text-to-Image Generation
@@
Recent research utilizing diffusion models have made significant strides in the realm of image generation. Notably, [14] has outperformed the majority of generative adversarial networks (GANs) in terms of generation metrics, indicating a substantial progression in the field. The use of diffusion models in the generation of images from text has produced remarkable outcomes.
@@
To align generated images and prompts, Glide [1] integrated diffusion models with guidance techniques. Large-scale diffusion models have been created [2, 3], exhibiting even more remarkable performance in generating images from text. However, the substantial size of these models renders them unsuitable for average users with limited computing capabilities. [4] proposed the latent diffusion model (LDM), which employs an autoencoder to map images from the image space to the latent space, thereby reducing the computational demands of diffusion models. Furthermore, LDM incorporates a cross-attention layer within its structure, achieving outstanding results in the text-to-image task. Our work builds on LDM, specifically utilizing a pretrained LDM called stable diffusion [4].
2.2 Image Editing
2.2.1 Text Driven Image Editing.
@@ 2.2 Image Editing
Recently, diffusion model-based image editing has achieved unparalleled results. For example, P2P [5] utilizes cross-attention control to edit images from one prompt to another, while PnP [6] introduces a plug-and-play method for text-driven image translation. Another training-free text-driven image translation technique is presented by pix2pix-zero [7], who use regularized forward DDIM to improve the editability of the inverted latent code and apply cross-attention guidance to maintain the image structure. PTI [8] edits specific objects, which involves updating the condition embedding during the image reconstruction process in stage one. The updated condition embedding is then employed in stage two to finalize the image editing. NTI [15] discovers that optimizing Null-text Embedding can accurately reconstruct the original image while avoiding fine-tuning the model and text embedding, thereby preserving the structure of image and enabling high-fidelity editing of real images.
2.2.2 Consistent Image Editing.
@@
Consistent image editing [11, 9] seeks to modify a single object in an image while leaving the rest of the image intact. DragGAN [16] was the first to introduce drag-style image editing to achieve this consistency. However, the quality and generalizability of DragGAN’s editing are constrained by its dependence on GANs. Capitalizing on the success of diffusion models, [12, 17] have employed diffusion models to execute drag-style image editing.
@@
Imagic [11] necessitates optimizing both textual embedding and diffusion model. It linearly interpolates the target text embedding and the optimized embedding, and then inputs the fine-tuned model to generate the edited image. MasaCtrl [9] employs mutual self-attention which involves querying image content from source images using a self-attention mechanism to achieve consistent image editing. These methods are restricted to modifying the actions of objects and lack control over the generated position. In contrast, ours consists of an inversion branch and an edit branch, allowing for simultaneous editing of the object’s action and control over its generated position.
Figure 2: Move&Act pipeline. In inversion stage, we update latent code and transfer object information to the target while repairing source background. In reverse stage, we use self-attention image features for consistent image editing.
3 Methodology
@@
3.1 Object Information Transfer
@@
3.1.1 Object Information Infusion.
Our observation reveals that the attention map, associated with each object token within the cross-attention mechanism of stable diffusion, encompasses spatial position information about the object within the image. This observation aligns with the findings of previous studies [18, 19]. Leveraging this insight, several studies [19, 20, 21, 22] have utilized this mechanism to manipulate the arrangement of objects within images. In a similar vein to [18, 19], we utilize a 16$\times$16 resolution attention map to update the latent code, as this particular resolution attention map is rich in semantic information. Consequently, to facilitate the relocation of the object to the target area, we ensure that the attention map, corresponding to the edited object, achieves the highest attention score within the confines of the target area:
@@
3.1.2 Source Area Inpainting.
@@
In order to harmonize the background of the source area with that of the edge area, we extract features from both areas of the latent code and subsequently compute the $L_{1}$ loss between the features of these regions. DIFT [23] discovered that the UNet of the diffusion model is an efficient feature extractor, thus we utilize UNet to extract features from the latent code. The decoder of the UNet denoiser comprises four blocks of varying scales. According to DIFT, the second layer contains more semantic information, while the third layer encompasses more geometric information. In our method, we opt for the third layer.
@@
Apart from its impact on the source area, $L_{sai}$ also serves a potentially crucial role. Employing solely $L_{oii}$ may not facilitate the complete transfer of object information to the target area, even though $L_{oii}$ minimizes the attention scores of the non-target area. In Figure 4, when only $L_{oii}$ is utilized, remnants of the object information persist in the source area, which could hinder the generation of the edited object within the target area. By concurrently employing both $L_{sai}$ and $L_{oii}$, $L_{sai}$ well erases the object in the source area.
@@ 3.2 Background Preservation
$z_{t}=z_{t}-\alpha\cdot\nabla_{z_{t}}L_{inv}.$ (7)
3.3 Content Editing Consistency
In contrast to the approach presented in [9], we query the key and value during the inversion stage, as we believe that the key and value from this stage contain richer information about the original image’s object appearance and background compared to using the key and value from the reconstruction stage as in [9]. Furthermore, utilizing the key and value from the inversion stage diminishes the image reconstruction cost. Consequently, our optimization process transforms the original three branches of inversion, reconstruction, and editing [9] into two branches: inversion and editing. The exchange of key and value in the optimized operation can be expressed:
4 Experiments
@@
(i) (e.1) Input
(j)  (e.2) A sitting cat
(k) (f.1) Input
(l)  (f.2) A running cat
Figure 6: Visualization of our proposed Move&Act for image editing.
@@
4.1 Implementation Details
In the experimental section, we employ the stable diffusion [4] using checkpoint v1.4. We utilize a 50-step DDIM inversion process, and the classifier-free guidance is set at 7.5. During the 35-$th$ step of inversion, we update the latent code. The hyperparameter values for $\lambda_{oii}$, $\lambda_{sai}$, and $\lambda_{bg}$ in Eq. (6) are set to 0.5, 0.25, and 0.25, respectively. In order to update the latent code in accordance with Eq. (7), we establish the step size $\alpha$ at 150 and designate the number of iterations as 50. From the 7-$th$ step of the reverse process onward, we substitute the key and value within the self attention mechanism as per Eq (8). All experiments are conducted on a single 3090 GPU, well demonstrating the democratization of our introduced method in this paper.
4.2 Evaluation Dataset
@@ 4.3 Qualitative Result
We configure the generated position of the edited object to match the object’s original location, for comparing with existing methods. As shown in Figure 1 and Figure 7, our method exhibits superior background preservation, and the actions of the edited objects align more closely with the user-provided prompts. This comparison highlights our method in achieving enhanced consistency.
4.4 Quantitative Result
@@
Method CLIP-Score $AP_{50}$
InstructPix2Pix 0.2966 -
PnP 0.3034 -
MasaCtrl 0.3068 -
Move&Act (Ours) 0.3101 32.9
We calculate the following metrics to compare with previous methods: (1) CLIP Score [24], which assesses the association between text and images. A higher CLIP Score value signifies a greater correlation between text-image pairs. (2) YOLO Score [25], which evaluates the accuracy of the generated position of the edited object using the $AP_{50}$ metric. To detect the bounding box and predict the category of the edited object in the generated image, we employ YOLOv8 [26].
@@
(e) Input
(f) Step 25
(g) Step 35
(h) Step 45
@@
4.5 Ablation Study
4.5.1 Time Step to Update $z_{t}$ During Inversion.
@@ 4.5 Ablation Study
4.5.2 Iterations of Latent Updating.
Determining the appropriate number of iterations for converging Eq. (6) is essential for achieving optimal results. We conduct experiments with varying iteration settings, as demonstrated in Figure 9. When the iteration number is set to 25, the edited image struggles to maintain the original background effectively, and the cat’s action appears distorted. Conversely, setting the iteration number to a higher value of 75 allows for better preservation of the original background; however, the cat’s action does not align well with the user’s provided prompt.
@@
4.5.3 Starting Time Step for Editing Consistency.
4.5.4 Self-attention Layer Setting.
5 Conclusion
@@ Acknowledgement
This work was supported by National Science and Technology Major Project (No. 2022ZD0118201), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of Chine a (No. U21B2037, No. U22B2051, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001).

### v2

@@ Abstract
(e) Original image
(g) (g) PnP
(h) (e) Input
(j) (i)   Move&Act
Figure 1: (a) - (d): our method allows users to control where to place the edited object. (e) - (i): altering only action, our method captures the “standing” pose and maintains consistent backgrounds with the input image.
(f) (f) InstructPix2Pix
(i) (h) MasaCtrl
Introduction
Recent advancements in text-to-image generation using diffusion models (Nichol et al. 2022; Saharia et al. 2022; Ramesh et al. 2022; Rombach et al. 2022) have significantly improved image generation capabilities. By capitalizing on extensive training data, these models (Saharia et al. 2022; Ramesh et al. 2022) excel in generating remarkable images based on user prompts. Compared to generative adversarial networks (GANs), diffusion models provide superior quality and diversity in image generation. Consequently, numerous studies (Hertz et al. 2023; Tumanyan et al. 2023; Parmar et al. 2023; Dong et al. 2023; Cao et al. 2023; Huang et al. 2023) have begun investigating diffusion models for image editing tasks. Predominantly, these methods involve text-conditioned editing, where users supply text input to modify images. Although techniques like (Hertz et al. 2023; Tumanyan et al. 2023; Parmar et al. 2023; Dong et al. 2023) can adeptly edit image style, object appearance, and object category based on text input, they do not alter scene layout or modify object actions in the image. Nonetheless, there are instances where users desire to edit an object’s action while maintaining its appearance and background. This specific editing process is called consistent image editing.
Figure 2: Move&Act pipeline. In the inversion stage, we update latent code and transfer object information to the target area while repairing the source background. In the reverse stage, we use self-attention image features for consistent image editing.
Imagic (Kawar et al. 2023) adopts a complex prompt fine-tuning process for text-based consistent image editing. MasaCtrl (Cao et al. 2023) offers a tuning-free approach, transforming self-attention into mutual self-attention. These methods edit an object’s action without allowing control over the object’s generated position. DragonDiffusion (Mou et al. 2024) facilitates object displacement within the image, yet it restricts the editing of the relocated object’s action. In practical applications, users may require simultaneous editing of an object’s action and position. As illustrated in the upper part of Figure 1d, we allow users to input editing text to adjust the object’s action, while also providing a bounding box to dictate the object’s generated position.
We also maintain better background preservation. By designating the object’s generation position as its original location, our method effectively reduces to MasaCtrl’s editing function, solely modifying the object’s action. In the lower part of Figure 1d, the backgrounds generated by InstructPix2Pix (Brooks, Holynski, and Efros 2023), PnP (Tumanyan et al. 2023) and MasaCtrl (Cao et al. 2023) exhibit a certain loss in detail, whereas our method produces a background more closely resembling the input image. Also, the cat’s action in our results aligns more accurately with the editing prompt and the surrounding environment.
@@ Introduction
Related work
Text-to-Image Generation
@@
Recent research utilizing diffusion models have made significant strides in the realm of image generation. Notably, (Dhariwal and Nichol 2021) has outperformed the majority of generative adversarial networks (GANs) in terms of generation metrics, indicating a substantial progression in the field. The use of diffusion models in the generation of images from text has produced remarkable outcomes.
@@
To align generated images and prompts, Glide (Nichol et al. 2022) integrated diffusion models with guidance techniques. Large-scale diffusion models have been created (Saharia et al. 2022; Ramesh et al. 2022), exhibiting even more remarkable performance in generating images from text. However, the substantial size of these models renders them unsuitable for average users with limited computing capabilities. (Rombach et al. 2022) proposed the latent diffusion model (LDM), which employs an autoencoder to map images from the image space to the latent space, thereby reducing the computational demands of diffusion models. Furthermore, LDM incorporates a cross-attention layer within its structure, achieving outstanding results in the text-to-image task. Our work builds on LDM, specifically utilizing a pretrained LDM called stable diffusion (Rombach et al. 2022).
Image Editing
Text Driven Image Editing.
@@
Recently, diffusion model-based image editing has achieved unparalleled results. For example, P2P (Hertz et al. 2023) utilizes cross-attention control to edit images from one prompt to another, while PnP (Tumanyan et al. 2023) introduces a plug-and-play method for text-driven image translation. Pix2Pix-Zero (Parmar et al. 2023) utilizes regularized forward DDIM and cross-attention guidance to achieve text-driven image translation. PTI (Dong et al. 2023) edits particular objects by optimizing the condition embedding during the image reconstruction process. This updated condition embedding is then utilized to finalize the image editing. NTI (Mokady et al. 2023) discovers that optimizing null-text Embedding can accurately reconstruct the original image while avoiding fine-tuning the model, thereby preserving the structure of image and enabling high-fidelity editing of real images.
Consistent Image Editing.
@@
Consistent image editing (Kawar et al. 2023; Cao et al. 2023) seeks to modify a single object in an image while leaving the rest of the image intact. DragGAN (Pan et al. 2023) was the first to introduce drag-style image editing to achieve this consistency. However, the quality and generalizability of DragGAN’s editing are constrained by its dependence on GANs. Capitalizing on the success of diffusion models, (Mou et al. 2024; Shi et al. 2023) have employed diffusion models to execute drag-style image editing, resulting in promising outcomes.
@@
Imagic (Kawar et al. 2023) necessitates optimizing both textual embedding and diffusion model. It linearly interpolates the target text embedding and the optimized embedding, and then inputs the fine-tuned model to generate the edited image. MasaCtrl (Cao et al. 2023) employs mutual self-attention which involves querying image content from source images using a self-attention mechanism to achieve consistent image editing. These methods are restricted to modifying the actions of objects and lack control over the generated position. In contrast, ours consists of an inversion branch and an edit branch, allowing for simultaneous editing of the object’s action and control over its generated position.
Methodology
@@
Object Information Transfer
@@
Object Information Infusion.
Our observation reveals that the attention map, associated with each object token within the cross-attention mechanism of stable diffusion, encompasses spatial position information about the object within the image. This observation aligns with the findings of previous studies (Chefer et al. 2023; Xie et al. 2023). Leveraging this insight, several studies (Xie et al. 2023; Xue et al. 2023; Chen, Laina, and Vedaldi 2024; Phung, Ge, and Huang 2023) have utilized this mechanism to manipulate the arrangement of objects within images. In a similar vein to (Chefer et al. 2023; Xie et al. 2023), we utilize a 16$\times$16 resolution attention map to update the latent code, as this particular resolution attention map is rich in semantic information. Consequently, to facilitate the relocation of the object to the target area, we ensure that the attention map, corresponding to the edited object, achieves the highest attention score within the confines of the target area:
@@
Source Area Inpainting.
@@
In order to harmonize the background of the source area with that of the edge area, we extract features from both areas of the latent code and subsequently compute the $L_{1}$ loss between the features of these regions. DIFT (Tang et al. 2023) discovered that the UNet of the diffusion model is an efficient feature extractor, thus we utilize UNet to extract features from the latent code. The decoder of the UNet denoiser comprises four blocks of varying scales. According to DIFT, the second layer contains more semantic information, while the third layer encompasses more geometric information. In our method, we opt for the third layer.
@@
Apart from its impact on the source area, $L_{sai}$ also serves a potentially crucial role. Employing solely $L_{oii}$ may not facilitate the complete transfer of object information to the target area, even though $L_{oii}$ minimizes the attention scores of the non-target area. In Figure 4, when only $L_{oii}$ is utilized, remnants of the object information persist in the source area, which could hinder the generation of the edited object within the target area. By concurrently employing $L_{sai}$ and $L_{oii}$, $L_{sai}$ will completely erase the object in the source area.
@@
Figure 6: Visualization of our proposed Move&Act for consistent image editing.
$z_{t}=z_{t}-\alpha\cdot\nabla_{z_{t}}L_{inv}.$ (7)
Content Editing Consistency
In contrast to the approach presented in (Cao et al. 2023), we query the key and value during the inversion stage, as we believe that the key and value from this stage contain richer information about the original image’s object appearance and background compared to using the key and value from the reconstruction stage as in (Cao et al. 2023). Furthermore, utilizing the key and value from the inversion stage diminishes the image reconstruction cost. Consequently, our optimization process transforms the original three branches of inversion, reconstruction, and editing (Cao et al. 2023) into two branches: inversion and editing. The exchange of key and value in the optimized operation can be expressed:
Experiments
@@
Implementation Details
In the experimental section, we employ the stable diffusion (Rombach et al. 2022) using checkpoint v1.4. We utilize a 50-step DDIM inversion process, and the classifier-free guidance is set at 7.5. During the 35-$th$ step of inversion, we update the latent code. The hyperparameter values for $\lambda_{oii}$, $\lambda_{sai}$, and $\lambda_{bg}$ in Eq. (6) are set to 0.5, 0.25, and 0.25, respectively. In order to update the latent code in accordance with Eq. (7), we establish the step size $\alpha$ at 150 and designate the number of iterations as 50. From the 7-$th$ step of the reverse process onward, we substitute the key and value within the self attention mechanism as per Eq (8). All experiments are conducted on a single 3090 GPU, well demonstrating the democratization of our introduced method in this paper.
Evaluation Dataset
@@
We configure the generated position of the edited object to match the object’s original location, for comparing with existing methods. As shown in Figure 1d and Figure 7, our method exhibits superior background preservation, and the actions of the edited objects align more closely with the user-provided prompts. This comparison highlights our method in achieving enhanced consistency.
Quantitative Result
We calculate the following metrics to compare with previous methods: (1) CLIP Score (Radford et al. 2021), which assesses the association between text and images. A higher CLIP Score value signifies a greater correlation between text-image pairs. (2) YOLO Score (Li et al. 2021), which evaluates the accuracy of the generated position of the edited object using the $AP_{50}$ metric. To detect the bounding box and predict the category of the edited object in the generated image, we employ YOLOv8 (Reis et al. 2023).
Method CLIP-Score $AP_{50}$
InstructPix2Pix 0.2966 $\pm$ 0.0013 -
PnP 0.3035 $\pm$ 0.0009 -
MasaCtrl 0.3066 $\pm$ 0.0008 -
Move&Act (Ours) 0.3098 $\pm$ 0.0011 33.70 $\pm$ 0.74
@@
(e) (a) Input
(f) (b) Step 25
(g) (c) Step 35
(h) (d) Step 45
@@
Ablation Study
Time Step to Update $z_{t}$ During Inversion.
@@
Iterations of Latent Updating.
We perform experiments to find the suitable number of iterations to converge Eq. (6), as shown in Figure 9. With 25 iterations, the edited image struggles to maintain the original background effectively, and the cat’s action appears distorted. Conversely, setting the iteration number to a higher value of 75 allows for better preservation of the original background; however, the cat’s action does not align well with the user’s provided prompt.
@@
Starting Time Step for Editing Consistency.
Self-attention Layer Setting.
@@
Limitation and Future Work
Complex image input is a challenge for our method, particularly in effectively restoring the source area’s background when it is surrounded by complex background. Additionally, the kernel size choice for the dilation operator is pivotal in our approach. A smaller kernel size results in an edge area with deficient features, leading to suboptimal outcomes. Conversely, an excessively large kernel size produces a broad edge area, encompassing background regions too distant from the source area to be relevant for restoration. To refine our method, future research will concentrate on investigating source area inpainting loss with an enhanced capacity for background recovery, thereby ensuring more precise and consistent results across various editing scenarios.
Conclusion
@@ Acknowledgement
This work was supported by National Science and Technology Major Project (No. 2022ZD0118201), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. U23A20383, No. U21A20472, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No. 2021J06003, No.2022J06001).